The equilibrium level is R17 000m. The full employment income is R22 000m and the marginal propensity to consume is 0,8. By how much should the investment expenditure change to bring nations income to full employment equilibrium?
1
Expert's answer
2020-04-14T10:03:02-0400

equilibrium level, E = R17000m

 full employment income,Fe =R22000m

"\text{change in income}=Fe-E"

change in income=R22000m-R17000m=R5000m

hence the income expenditure must change by;

"R5000m\times 0.8" =R4000m